# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

### v3

@@ Abstract
Online multi-object tracking (MOT) plays a pivotal role in autonomous systems. The state-of-the-art approaches usually employ a tracking-by-detection method, and data association plays a critical role. This paper proposes a learning and graph-optimized (LEGO) modular tracker to improve data association performance in the existing literature. The proposed LEGO tracker integrates graph optimization and self-attention mechanisms, which efficiently formulate the association score map, facilitating the accurate and efficient matching of objects across time frames. To further enhance the state update process, the Kalman filter is added to ensure consistent tracking by incorporating temporal coherence in the object states. Our proposed method utilizing LiDAR alone has shown exceptional performance compared to other online tracking approaches, including LiDAR-based and LiDAR-camera fusion-based methods. LEGO ranked $1^{st}$ at the time of submitting results to KITTI object tracking evaluation ranking board and remains $2^{nd}$ at the time of submitting this paper, among all online trackers in the KITTI MOT benchmark for cars [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking.php.
@@ I Introduction
Multi-object tracking (MOT) is a crucial technology utilized in various systems within the field of intelligent transportation systems (ITS), including advanced driver assistance systems (ADAS) and autonomous driving (AD). Fig. 1 illustrates the task of LiDAR-based MOT, which presents several challenges that need to be addressed:
@@
To tackle the aforementioned challenges, an appealing approach is by redefining the data association problem as a bipartite graph matching problem, which takes advantage of its graphical formulation. In this redefined context, a self-attention-based GNN is integrated with the Kalman filter to facilitate data association and perform state updates.
Our approach incorporates a self-attention-based GNN optimization module, which operates by reasoning over non-Euclidean graph data through message passing. This module enhances the relational reasoning abilities of the system and reduces information redundancy. By leveraging self-attention mechanisms, our GNN can effectively capture long-range dependencies and complex relationships between objects, contributing to improved data association accuracy. Additionally, an offset module is specifically designed to refine the estimated bounding boxes from the detector. The primary objective of this module is to correct the output of the detector, resulting in more accurate and reliable information about the object’s motion.
@@
  • A learning and graph-optimized (LEGO) modular tracker is proposed for online MOT. The tracker is built upon recent model-based tracker, i.e. PC3T [38], with introducing two additional learning modules, an offset correction module and a self-attention based GNN, which are easily comprehensible, tunable, and adaptable to different sensor modalities.
  • The offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities. The graph dual attention network (GDAN) is introduced to process the bipartite graph input to reduce the information redundancy, and directly generate the association score matrix.
  • Empirical evaluation and performance analysis: The paper empirically evaluates the proposed GDAN module, demonstrating its effectiveness in enhancing tracking performance. The proposed LEGO is evaluated using the KITTI MOT benchmark, and the results showcase commendable performance in terms of tracking accuracy and reliability.
@@ II-A1 Model-based Methods
There are alternative algorithms where each potential object is modelled using a Bernoulli process with probabilistic object existence. RFS-M3 [18] utilized the Poisson multi-Bernoulli mixture filter based on a random finite set (RFS) to tackle the LiDAR-based MOT problem. Through systematic comparative analysis, GNN-PMB [28] demonstrated that the contemporary RFS-based Bayesian tracking framework outperformed the traditional random vector-based Bayesian tracking framework. BP-Tracker [24] presented a factor graph formulation of the MOT problem and employed a belief propagation algorithm to compute the marginal association probability, representing a significant advancement in the field.
@@ II-B MOT with LiDAR and Camera Fusion
The effectiveness of MOT can be further enhanced by adopting a fusion approach that combines data from LiDAR and camera sources. Several methods, such as Probabilistic3DMM [40], CBMOT, GNN3DMOT [43], and AlphaTrack [42] have utilized this fusion approach to leverage the complementary strengths of different sensor modalities, leading to more comprehensive and precise object tracking. IMSF MOT [44] proposes a novel feature fusion method using Pointnet++ to extract more discriminative features and improve multiple object tracking performance.
DeepFusionMOT [48] and EagerMOT [45] improve tracking performance compared to Probabilistic3DMM by implementing a two-stage data association scheme. This scheme leverages 3D detection data obtained from LiDAR and camera inputs, and 2D detection data obtained solely from the camera. By combining information from multiple modalities, a more comprehensive understanding of the environment is achieved, potentially enhancing the robustness and accuracy of the tracking process.
Building upon EagerMOT, AlphaTrack introduces a feature extractor that concatenates image and point cloud information as input to enhance performance. Additionally, other relevant works, such as DualTracker [47], HIDMOT [50], CAMO-MOT [52], MSA-MOT [51], and JMODT [46] utilize PointGNN or PointRCNN as their 3D object detectors to estimate object detections. They are combined with detection proposals from image data, and a hybrid multi-modal input is employed for the association within the tracking-by-detection task, potentially improving the overall performance and accuracy of the tracking system.
@@ III The Proposed LEGO Modular Tracker
Figure 2: The proposed LEGO tracker seamlessly integrates several modules. First, the offset correction module refines detection results and corrects detected errors. Simultaneously, the motion prediction module utilizes the tracked state $\bm{X}_{t}$ to forecast the next state $\bm{X}_{t|t-1}$. The feature extraction module then derives 3D feature maps for both predicted and detected objects. Using this information, the GDAN score calculation module constructs a bipartite graph and computes the GDAN score matrix. The association module subsequently integrates the geometry and motion cost matrix with the GDAN score matrix, yielding an association map. Finally, the state update module refines the predicted state, and the track management module processes the matched detected states, completing the tracking cycle.
Within this section, we will introduce the LEGO modular tracker. Firstly, we will present its overall framework, followed by a detailed elaboration of each module. Fig. 2 shows the proposed LEGO tracker, which comprises several modules: motion prediction, offset correction, feature extraction, GDAN score calculation, association, state update, and track management.
@@ III-A The Framework of LEGO Modular Tracker
Initiating the process, the 3D point cloud is fed into a 3D object detector, which is responsible for identifying potential objects within the scene. To augment the accuracy of the detection results, an offset correction module is proposed. The primary function of the offset correction module is to anticipate the offset between the detection result and the ground truth, thereby smoothing the detector’s outputs. Meanwhile, the motion prediction module takes the tracked state from the previous timestep and predicts the state of each object in the current timestep by Kalman prediction. Then, the geometry and motion cost matrix $\bm{A}$ is calculated by the detected state (after the offset correction module) and the predicted state from the motion prediction module.
@@
Upon inputting the bipartite graph $\bm{G}$ and the init GDAN score matrix into the GDAN module, the GDAN score matrix $\bm{B}$ is produced as an output. The final association cost matrix is then obtained via the proposed method, encapsulating both the GDAN score matrix and the geometry and motion cost matrix. The data association module is utilized to discern the matched pairs. Finally, the association map is fed into the state update module, prompting an update in the detected state within the aforementioned module being conveyed to the track management module.
@@ III-B Motion Prediction Module
Recently, LSTM networks have become increasingly popular in MOT methods for predicting changes in the object state [57]. These networks are favoured for their ability to handle sequential data exceptionally well. However, it is important to note that LSTM networks require more time than traditional motion prediction techniques, like those based on the Kalman filter. Certain MOT techniques prefer using a constant velocity motion model rather than LSTM networks. However, this model assumes that the object moves at a nearly constant speed, which may not hold true in several real-world situations. As a result, if there are consecutive missed detections, it could lead to considerable errors in motion prediction. To overcome this problem, the constant acceleration (CA) motion model [16] is used, which offers a more precise representation of the object state. In the CA motion model, the object state $\bm{X}_{t-1}$ at a specific temporal instance $t-1$ is represented as $\left[\bm{pos}_{t-1},\bm{v}_{t-1},\bm{\alpha}_{t-1}\right]^{T}$, where $\bm{pos}_{t-1}$ denotes the position, $\bm{v}_{t-1}$ represents the velocity, and $\bm{\alpha}_{t-1}$ indicates the orientation of the object. The prediction of the mean state $\bm{X}_{t|t-1}$ and covariance $\bm{P}_{t|t-1}$ using Kalman prediction equations with the CA motion model is as follows:
@@ III-D Feature Extraction Module
Figure 4: The structure of the feature extraction module. The 3D point cloud is served as input into the transformation network (T-Net), where they undergo augmentation before being fed into the PointNet architecture. Following this, the outputs of the penultimate layer of the PointNet are channelled into the multi-layer perception (MLP). This subsequent step yields the desired outputs: the prediction node feature and the detection node feature.
@@ III-D2 Feature Embedding
Sequentially, the 3D point cloud feature map is fed into the MLP-based embedding network, to extract embedded features as illustrated in Fig. 4. This process can be described as
@@
where $\bm{F}$ denotes the output feature maps of the feature extraction module.
III-E GDAN Score Calculation Module
@@
Utilizing the feature map extracted from the feature extraction module, a GDAN score matrix, denoted as $\bm{B}$, is calculated through the GDAN score calculation module. Similar to [40], the purpose of employing such matrix is to extract the similarities between different detections and predicted tracks. This calculation takes into consideration the relationships between point cloud voxels of various objects. We first construct the graph. Most of the recent works which applied GNN for 3D MOT, e.g. PTP [53], Rethinking3DMOT [37] , 3Dmotformer [54], etc, use the fully connected graph with dimension of $\left(J+U\right)\times\left(J+U\right)$. However, such graph entails redundant computations due to the simultaneous inclusion of detection and prediction nodes. To this end, we follow GNN3DMOT [43] to use the prior knowledge that the matching should only happen across frames that we construct a bipartite graph. The structure of the graph is represented as a matrix of dimensions $J\times U$, where $J$ represents the number of detection nodes and $U$ denotes the number of prediction nodes. The bipartite graph mitigate this computational redundancy, effectively reducing the dimension of the association score matrix to $J\times U$. Consequently, the association between predicted and detected objects is framed as a maximum-weighted bipartite matching problem, wherein each edge in the graph corresponds to an association score between two distinct nodes. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with $\bm{N}$ consisting of two node sets, $\bm{N}_{det}$ (the set of nodes representing detected objects) and $\bm{N}_{pre}$ (the set of nodes representing predicted objects). The Cardinalities of these sets are $\left|\bm{N}_{det}\right|$ = $J$ and $\left|\mathbf{N}_{pre}\right|=U$, respectively. The set of edges, $\bm{E}$, captures the associations between nodes, and each edge within $\bm{E}$ encodes an association score representing the strength of the relationship between the corresponding nodes. To initialize the bipartite graph, all edge values are assigned an initial score of 1.
Figure 5: The architecture of GDAN module. In the initial stage, a bipartite graph is fed into the multi-head self-attention (MSA) block. This data together with the normalized initialization GDAN association score matrix in the combined framework of message passing and update block. This block includes several components: an association self-attention mechanism that adjusts the original self-attention, a normalization layer, and an MLP. Finally, this comprehensive system yields the output in the form of an association score matrix.
Subsequently, the initial GDAN score matrix denoted as $\bm{B}_{init}$, in conjunction with the predicted feature map and detected feature map, serves as input parameters for further processing. This sequential computation culminates in the generation of a GDAN score matrix denoted as $\bm{B}\in\mathbb{R}^{J\times U}$, representing the importance of each pairwise similarity in the final scoring scheme.
@@
The method in [59] directly applies polynomials to update the feature information on the node through message passing. This could result in redundant information for each node since different nodes should contribute different useful information instead of having the same contribution for message passing. To overcome the limitations in [59], GNN3DMOT [43] use graph attention network (GAT) [60] to realize the message passing and updating. However, GAT belongs to the spatial convolution and process the message passing and updating node by node which ignoring the whole structure of the graph. This could miss the important information in the graph we construct. To overcome these limitations, we proposed the GDAN module. We first use an MSA encoder to simple extract the feature representation of each node. Considering the input are point cloud in two sequence frames, we also design an embedding network with LSTM and several fully connect layers to further extract the embedding representation of features, then a modified self-attention-based mechanism is incorporated to manage message passing and updating within the bipartite graph. Different from directly using self-attention like GAT, we add a learnable adjacency matrix into the calculation of the self-attention during the message passing and updating process which can provide the structure information. Our proposed modified self-attention blocks are specifically tailored to suit the bipartite graph structure, as depicted in Fig. 5.
The association self-attention calculation method undergoes modification to enable the learning of association scores via:
$Attention\left(\bm{Q},\bm{K},\bm{V},\bm{B}\right)=f\left(\bm{Q},\bm{K},\bm{V}\right)\times\bm{W}\times\bm{B}^{T},$ (5)
$f\left(\bm{Q},\bm{K},\bm{V}\right)=softmax\left(\frac{\bm{QK^{T}}}{\sqrt{d_{k}}}\right)\times\bm{V},$ (6)
where $\bm{Q}$, $\bm{K}$, and $\bm{V}$ denote the query, key, and value obtained from the feature maps, respectively. This method enables varying weights to be applied to the information provided by different nodes to each node in the graph. By assigning attention values to each node, data from various nodes can be combined to update the edge score, which helps to avoid duplication and loss of information. Additionally, the GDAN score matrix can be obtained directly from the network without changing its size. To generate the GDAN score matrix $\bm{B}$, a pre-learning phase is conducted using the self-attention block. The entry $B_{j,u}$ at the $j$-th row and $u$-th column of matrix $\bm{B}$ is
$B_{j,u}=MLP\left(Attention\left(\bm{Q},\bm{K},\bm{V},\bm{B}_{init}\right)\right).$ (7)
@@
The corresponding loss function for training the network to generate GDAN score matrix $\bm{B}$, association matrix loss, is expressed mathematically as
@@
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware $y_{j,u}$ is a binary indicator which is either 00 or $1$ as it represents the ground truth of the match between the $u$-th predicted object and the $j$-th detected object. When there is a clear match, $y_{j,u}$ is assigned as 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, we define a match based on object identity continuity between time frames: a pair is considered as ”matched” if the closest ground-truth box to the tracking box in the previous time frame and the closest ground-truth box to the detection box in the current time frame have the same index and their IoU with their closest ground-truth boxes are more than 0.7. The value of $B_{j,u}$ is obtained from the GDAN score matrix $\bm{B}$ and ranges between 0 and 1.

### v4

@@ Abstract
Online multi-object tracking (MOT) plays a pivotal role in autonomous systems. The state-of-the-art approaches usually employ a tracking-by-detection method, and data association plays a critical role. This paper proposes a learning and graph-optimized (LEGO) modular tracker to improve data association performance in the existing literature. The proposed LEGO tracker integrates graph optimization, which efficiently formulate the association score map, facilitating the accurate and efficient matching of objects across time frames. To further enhance the state update process, the Kalman filter is added to ensure consistent tracking by incorporating temporal coherence in the object states. Our proposed method utilizing LiDAR alone has shown exceptional performance compared to other online tracking approaches, including LiDAR-based and LiDAR-camera fusion-based methods. LEGO ranked $3^{rd}$ among all trackers (both online and offline) and $2^{nd}$ among all online trackers in the KITTI MOT benchmark for cars [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking.php, at the time of submitting results to KITTI object tracking evaluation ranking board.
@@ I Introduction
Multi-object tracking (MOT) is a crucial technology utilized in various systems within the field of intelligent transportation systems (ITS), like traffic monitoring [1, 2, 3], advanced driver assistance systems (ADAS) [4] and autonomous driving (AD) [5]. Fig. 1 illustrates the task of LiDAR-based MOT, which presents several challenges that need to be addressed:
@@
To tackle the aforementioned challenges, an appealing approach is by redefining the data association problem as a bipartite graph matching problem, which takes advantage of its graphical formulation. In this redefined context, a self-attention-based GNN combined with graph structure information is integrated with the Kalman filter to facilitate data association and perform state updates.
Our approach incorporates a new GNN optimization module, which operates by reasoning over non-Euclidean graph data through message passing and combining the graph structure information. This module enhances the relational reasoning abilities of the system and reduces information redundancy. By leveraging self-attention mechanisms and graph structure information, our GNN can effectively capture long-range dependencies and complex relationships between objects, contributing to improved data association accuracy. Additionally, an offset module is specifically designed to refine the estimated bounding boxes from the detector. The primary objective of this module is to correct the output of the detector, resulting in more accurate and reliable information about the object’s motion.
@@
  • A learning and graph-optimized (LEGO) modular tracker is proposed for online MOT. Compared to previous model-based trackers, LEGO introduces two learning modules as its core, an offset correction module and a new GNN, which are easily comprehensible, tunable, and adaptable to different sensor modalities.
  • The offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities. The adjency matrix guide graph network (AMGN) is introduced to use the information of the graph structure to do the message passing and updating with self-attention, and directly generate the association score matrix.
  • Empirical evaluation and performance analysis: The paper empirically evaluates the proposed AMGN module, demonstrating its effectiveness in enhancing tracking performance. The proposed LEGO is evaluated using the KITTI MOT benchmark, and the results showcase commendable performance in terms of tracking accuracy and reliability.
@@ II-A1 Model-based Methods
As the raise of applying random finite set (RFS) in MOT applications [4], there are alternative algorithms where each potential object is modelled using a Bernoulli process with probabilistic object existence. RFS-M3 [23] utilized the Poisson multi-Bernoulli mixture filter based on a random finite set (RFS) to tackle the LiDAR-based MOT problem. Through systematic comparative analysis, GNN-PMB [33] demonstrated that the contemporary RFS-based Bayesian tracking framework outperformed the traditional random vector-based Bayesian tracking framework. BP-Tracker [29] presented a factor graph formulation of the MOT problem and employed a belief propagation algorithm to compute the marginal association probability, representing a significant advancement in the field.
@@ II-B MOT with LiDAR and Camera Fusion
The effectiveness of MOT can be further enhanced by adopting a fusion approach that combines data from LiDAR and camera sources. Several methods, such as Probabilistic3DMM [46], CBMOT [47], GNN3DMOT [50], and MF-Net [48] have utilized this fusion approach to leverage the complementary strengths of different sensor modalities, leading to more comprehensive and precise object tracking. IMSF MOT [52] proposes a novel feature fusion method using Pointnet++ to extract more discriminative features and improve multiple object tracking performance.
DeepFusionMOT [56] and EagerMOT [53] improve tracking performance compared to Probabilistic3DMM by implementing a two-stage data association scheme. This scheme leverages 3D detection data obtained from LiDAR and camera inputs, and 2D detection data obtained solely from the camera. By combining information from multiple modalities, a more comprehensive understanding of the environment is achieved, potentially enhancing the robustness and accuracy of the tracking process. Building upon EagerMOT, AlphaTrack [49] introduces a feature extractor that concatenates image and point cloud information as input to enhance performance. Additionally, other relevant works, such as DualTracker [55], HIDMOT [59], CAMO-MOT [5], MSA-MOT [60], and JMODT [54] utilize PointGNN or PointRCNN as their 3D object detectors to estimate object detections. They are combined with detection proposals from image data, and a hybrid multi-modal input is employed for the association within the tracking-by-detection task, potentially improving the overall performance and accuracy of the tracking system.
@@ III The Proposed LEGO Modular Tracker
Figure 2: The proposed LEGO tracker seamlessly integrates several modules. First, the offset correction module refines detection results and corrects detected errors. Simultaneously, the motion prediction module utilizes the tracked state $\bm{X}_{t}$ to forecast the next state $\bm{X}_{t|t-1}$. The feature extraction module then derives 3D feature maps for both predicted and detected objects. Using this information, the AMGN score calculation module constructs a bipartite graph and computes the AMGN score matrix. The association module subsequently integrates the geometry and motion cost matrix with the AMGN score matrix, yielding an association map. Finally, the state update module refines the predicted state, and the track management module processes the matched detected states, completing the tracking cycle.
Within this section, we will introduce the LEGO modular tracker. Firstly, we will present its overall framework, followed by a detailed elaboration of each module. Fig. 2 shows the proposed LEGO tracker, which comprises several modules: motion prediction, offset correction, feature extraction, AMGN score calculation, association, state update, and track management.
@@ III-A The Framework of LEGO Modular Tracker
Initiating the process, the 3D point cloud is fed into a 3D object detector, which is responsible for identifying potential objects within the scene. To augment the accuracy of the detection results, an offset correction module is proposed. The primary function of the offset correction module is to anticipate the offset between the detection result and the ground truth, thereby smoothing the detector’s outputs. Meanwhile, the classical motion prediction module takes the tracked state from the previous timestep and predicts the state of each object in the current timestep by Kalman prediction. Then, following the recent model-based offline tracker, i.e. PC3T [44], the geometry and motion cost matrix $\bm{A}$ is calculated by using the detected state after the offset correction module and the predicted state from the motion prediction module.
@@
Upon inputting the bipartite graph $\bm{G}$ and the learnable adjacency matrix into the proposed AMGN module, the AMGN score matrix $\bm{B}$ is produced as an output. The AMGN module employs a learnable adjacency matrix, which encapsulates the structural information of the graph. This design enables the model to progressively refine its comprehension of the graph’s architecture. Furthermore, it facilitates the quantification and learning of the relational strengths between any given pair of nodes. The final association cost matrix is then obtained via the proposed method, encapsulating both the AMGN score matrix and the geometry and motion cost matrix. The data association module is utilized to discern the matched pairs. Finally, same as all classical model-based trackers, the association map is fed into the state update module, prompting an update in the object states using Kalman filter, within the output being conveyed to the track management module.
@@ III-B Motion Prediction Module
Recently, LSTM networks have become increasingly popular in MOT methods for predicting changes in the object state [66]. These networks are favoured for their ability to handle sequential data exceptionally well. However, it is important to note that LSTM networks require more time than traditional motion prediction techniques, like those based on the Kalman filter. Certain MOT techniques prefer using a constant velocity motion model rather than LSTM networks. Yet, this model assumes that the object moves at a nearly constant speed, which may not hold true in several real-world situations. As a result, if there are consecutive missed detections, it could lead to considerable errors in motion prediction. To overcome this problem, the constant acceleration (CA) motion model [21] is used, which offers a more precise representation of the object state. In the CA motion model, the object state $\bm{X}_{t-1}$ at a specific temporal instance $t-1$ is represented as $\left[\bm{pos}_{t-1},\bm{v}_{t-1},\bm{\alpha}_{t-1}\right]^{T}$, where $\bm{pos}_{t-1}$ denotes the position, $\bm{v}_{t-1}$ represents the velocity, and $\bm{\alpha}_{t-1}$ indicates the orientation of the object. The prediction of the mean state $\bm{X}_{t|t-1}$ and covariance $\bm{P}_{t|t-1}$ using Kalman prediction equations with the CA motion model is as follows:
@@ III-D2 Feature Embedding
Sequentially, the 3D point cloud feature map is fed into the MLP-based embedding network, to extract embedded features. This process can be described as
@@
where $\bm{Z}^{0}$ denotes the output feature maps of the feature extraction module.
III-E AMGN Score Calculation Module
@@
Utilizing the feature map extracted from the feature extraction module, a AMGN score matrix, denoted as $\bm{B}$, is calculated through the AMGN score calculation module. Similar to [46], the purpose of employing such matrix is to extract the similarities between different detections and predicted tracks. This calculation takes into consideration the relationships between point cloud voxels of various objects. We first construct the graph. Most of the recent works which applied GNN for 3D MOT, e.g. PTP [62], Rethinking3DMOT [43] , 3Dmotformer [63], etc, use the fully connected graph with dimension of $\left(J+U\right)\times\left(J+U\right)$. However, such graph entails redundant computations due to the simultaneous inclusion of detection and prediction nodes. To this end, we follow GNN3DMOT [50] to use the prior knowledge that the matching should only happen across frames that we construct a bipartite graph. The structure of the graph is represented as a matrix of dimensions $J\times U$, where $J$ represents the number of detection nodes and $U$ denotes the number of prediction nodes. The bipartite graph mitigate this computational redundancy, effectively reducing the dimension of the association score matrix to $J\times U$. Consequently, the association between predicted and detected objects is framed as a maximum-weighted bipartite matching problem, wherein each edge in the graph corresponds to an association score between two distinct nodes. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with $\bm{N}$ consisting of two node sets, $\bm{N}_{det}$ (the set of nodes representing detected objects) and $\bm{N}_{pre}$ (the set of nodes representing predicted objects). The Cardinalities of these sets are $\left|\bm{N}_{det}\right|$ = $J$ and $\left|\mathbf{N}_{pre}\right|=U$, respectively. The set of edges, $\bm{E}$, captures the associations between nodes.
Figure 4: AMGN score calculation module. In the initial stage, a bipartite graph is fed into the multi-head self-attention (MHSA) block along with an LSTM block to extract features. This data, together with the initialization of the learnable adjacency matrix, are fed into the AMGN module. The AMGN module performs node message passing and updating using the self-attention coefficients derived from the learnable adjacency matrix. Subsequently, the learnable adjacency matrix is updated with the new feature representation obtained after the message passing and updating process. The new feature representation and updated learnable adjacency matrix are then fed into the next AMGN module. The final output is the AMGN score matrix $\bm{B}$, which is the latest learnable adjacency matrix from the last AMGN module.
@@
The method in [68] directly applies polynomials to update the feature information on the node through message passing. This could result in redundant information for each node since different nodes should contribute different useful information instead of having the same contribution for message passing. To overcome the limitations in [68], GNN3DMOT [50] uses Graph Attention Network (GAT) [69] to realize the message passing and updating. However, GAT processes the message passing and updating node by node, ignoring the information of entire graph structure. This could miss important information in the constructed graph. To this end, different from the approach directly using self-attention like GAT, we propose the AMGN module, which introduces a learnable adjacency matrix into the calculation of self-attention during the message passing and updating process. The learnable adjacency matrix which contains the structure information of the graph, is updated iteratively as part of the training process, allowing the model to continuously refine its understanding of the graph structure and adapt to the evolving feature representations of the nodes through the layers and learn the strength of associations between any pair of nodes, ensuring that the attention mechanism considers both the node features and the graph’s structural information with the earnable adjacency matrix which reflect the structure of the whole graph. To use the AMGN module, we first use a multi-head self-attention (MHSA) encoder to simply extract the feature representation of each node. Considering the input is point cloud in two sequence frames, we also design an embedding network with LSTM and several fully connected layers to further extract the embedding representation of features as shown:
$\mathbf{Z}^{0}=LSTM(MHSA(\mathbf{F})).$ (5)
Then, AMGN modules are used to manage message passing and updating within the bipartite graph to get the final AMGN score matrix as shown in Fig. 4. We totally have K AMGN modules cross together and we initialize the learnable adjacency matrix of the graph as $\bm{B}^{0}$, its element $B^{0}_{j,u}$ represents the association score between nodes $j$ in the detected objects node set and nodes $u$ in the predicted objects set and the feature representation of each node is initialized by the Eq .5 and $\bm{Z^{0}}=(\bm{z}^{0}_{j},\dots,\bm{z}^{0}_{J},\bm{z}^{0}_{u},\dots,\bm{z}^{0}_{U})$. We have used total $K$ AMGN modules and suppose vector $\overrightarrow{\bm{W\odot Z}^{k-1}}$ is a vector that contains feature of all node in $(k-1)$-th AMGN module and the $u$-th node features $w_{u}z^{k-1}_{u}$ and $j$-th node feature $w_{j}z^{k-1}_{j}\in\overrightarrow{\bm{W\odot Z}^{k-1}}$. We also set $\overrightarrow{\bm{W\odot Z}^{k}}$ in $k$-th AMGN module. In our module, taking $k$-th AMGN module as an example, following GAT, the coefficient which is used for message passing and updating between node $j$ and node $u$ is calculated as:
$e^{k}_{j,u}=LeakyReLU(a^{T}[w_{u}z^{k-1}_{u}||w_{j}z^{k-1}_{j}]),$ (6)
where $a$ is the learnable weight vector, ||||| | denotes concatenation, and $\mathbf{W}$ is a learnable weight matrix. Different from GAT, which directly normalizes $e^{k}_{j,u}$ and does message passing and updating of node features, we modify the attention coefficients by combining them with the updated learnable adjacency matrix $\bm{B}^{k-1}$ from previous AMGN module as shown:
$e^{’k}_{j,u}=e^{k}_{j,u}B^{k-1}_{j,u},$ (7)
where $B^{k-1}_{j,u}$ is the association score between node $j$ and node $u$. Then we normalize the modified attention coefficients with the softmax function as:
$\alpha^{k}_{j,u}=\frac{exp(e^{’k}_{j,u})}{\sum exp(e^{’k}_{j,u})},$ (8)
then we perform message passing and updating of node features as shown:
$z^{’k}_{j}=LeakyReLU(\sum a_{j,u}w_{u}{z^{k}_{u}}),$ (9)
we do the same for the $u$-th node to get the updated feature representation of node $u$ as $z^{’}_{u}$. Finally, we update value in the learnable adjacency matrix $\bm{B}^{k}$ as shown:
$B^{k}_{j,u}=LeakyReLU(\mathbf{b}^{T}[z^{’k}_{j}||z^{’k}_{u}]),$ (10)
where $\mathbf{b}^{T}$ represents the learnable weight vector at the $k$-th module. The output of the $k$-th AMGN module, which includes the learnable adjacency matrix $\bm{B}^{k}$ and the updated vector $\overrightarrow{\bm{W\odot Z}^{{}^{\prime}k}}$, serves as the input to the subsequent $(k+1)$-th AMGN module. Ultimately, the AMGN score matrix $\bm{B}$ is defined as $\bm{B}^{K}$, which is the output of the final AMGN module.
@@
The corresponding loss function for training the network to generate AMGN score matrix $\bm{B}$, association matrix loss, is expressed mathematically as
@@
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware $y_{j,u}$ is a binary indicator which is either 00 or $1$ as it represents the ground truth of the match between the $u$-th predicted object and the $j$-th detected object. When there is a clear match, $y_{j,u}$ is assigned as 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, we define a match based on object identity continuity between time frames: a pair is considered as ”matched” if the closest ground-truth box to the tracking box in the previous time frame and the closest ground-truth box to the detection box in the current time frame have the same index and their IoU with closest ground-truth boxes are more than 0.7. The value of $B_{j,u}$ is obtained from the AMGN score matrix $\bm{B}$ and ranges between 0 and 1.